OpenStudy (anonymous):

if a side is lengthened by 7 cm, the area becomes 225^2 what is the orinal length of a side of the original square?

OpenStudy (anonymous):

so let the original size be x than you have the equation (x+7)^2=225^2 so x+7=225 or x+7=-225 x=218 or x=-218 (but that is silly, a length cannot be -)

OpenStudy (anonymous):

-232 sorry :-)

OpenStudy (anonymous):

if a is the original side length, then a+7 is the new side (as long as only ONE side is lengthened) a*(a+7)=225^2 a^2+7a=225^2 a^2 +7a-225^2=0 use quadratic formula and get BOTH answers and check them.

OpenStudy (anonymous):

oh true, if only 1 side is lengthened

OpenStudy (anonymous):

cathie you answer is 232

OpenStudy (anonymous):

cant be 232, 232^2 is larger than 225^2

OpenStudy (anonymous):

opps sorry tutordiffeq your right!